QUESTION 6: P6-13A Kane Ltd. had a beginning inventory on January 1 of 250 units of product SXL at a cost of $160 per unit. During the year, purchases were as follows:

Units

Unit Cost

Total Cost

Mar. 15

700

$150

$105,000

July 20

500

  145

   72,500

Sept. 4

450

  135

   60,750

Dec.  2

100

  125

   12,500

Kane uses a periodic inventory system. At the end of the year, a physical inventory count determined that there were 200 units on hand.

Instructions

(a) Determine the cost of goods available for sale.

(b) Determine the cost of the ending inventory and the cost of the goods sold using (1) FIFO and (2) average cost. (Use unrounded numbers in your calculation of the average unit cost but round to the nearest cent for presentation purposes in your answer.)
Answer #1

Answer :

(a)

Cost of goods available for sale $290,750

(b)

Ending inventory Cost of goods sold
1 FIFO $26,000 $264,750
2 Average cost $29,075 $261,675

Working Note :

(a) Cost of goods available for sale

Particulars Units Unit cost Total
January 1 250 $160 $40,000
Mar 15 700 $150 $105,000
July 20 500 $145 $72,500
Sept 4 450 $135 $60,750
Dec. 2 100 $125 $12,500
Cost of goods available for sale 2000 $290,750

(b) FIFO Method :

Production units = 250 + 700 + 500 + 450 + 100 = 2000

Ending inventory units = 200

Sales units = 2000 - 200 = 1800

Cost of goods sold

Particulars Units Unit cost Total
January 1 250 $160 $40,000
Mar 15 700 $150 $105,000
July 20 500 $145 $72,500
Sept 4 350 $135 $47,250
1800 $264,750

Ending inventory = 100*$135 + 100*$125 = $260,000

Average cost Method

Unit cost Total cost / No. of units available for sale
$290,750/2000 = 145.375
Cost of goods sold 1800*145.375 =$261,675
Ending inventory 200*$145.375 =29,075